find sec beta if csc beta equal square root of 2
Answer by lwsshak3(11628)

find sec beta if csc beta equal square root of 2
use x for beta
secx=1/cosx
cscx=√2=1/sinx
sinx=1/√2
cosx=√(1-sin^2x)=√1-1/2=√1/2=1/√2
secx=1/cosx=√2